Use simplex method to maximize 𝑓 = 3π‘₯ + 5𝑦 + 4𝑧 subject to the conditions 2π‘₯ + 3𝑦 ≀ 18 2π‘₯ + 5𝑦 ≀ 10 3π‘₯ + 2𝑦 + 4𝑧 ≀ 15 and π‘₯, 𝑦, 𝑧 β‰₯ 0.


1
Expert's answer
2021-09-17T03:45:38-0400

One of the theorems of simplex method states that the solution of the linear problem exists at one of the edge points. In our case the region is bounded by 6 planes: "2x+3y=18,2x+5y=10," "3x+2y+4z=15" , "x,y,z=0". It is possible to find such solution in Maple via the commands:

with(Optimization):

Minimize(-3*x-5*y-4*z,{2*x+3*y<=18,2*x+5*y<=10,3*x+2*y+4*z<=15,x>=0,y>=0,z>=0});

Since it is the minimization problem, the function has the opposite sign.

It is at the intersection of planes "x=0,y=2,z=2.75". The value is "21"for the original maximization problem. Thus, the solution is at the point "(0,2,2.75)". The value is "21".

The manual method of finding solution is the following:

One of conditions from the task formulation is: "3x+2y+4z\leq15". Thus, "f\leq3y+15". Using the fact that "x\geq0" and inequality "2x+5y\leq10" we get: "y\leq2". Thus, "f\leq21". The point that satisfies equality "f=21" is: "x=0,y=2,z=2.75". It satisfies equalities "3x+2y+4z=15", "2x+5y=10" and "x=0".